### v1

# “Would You Want an AI Tutor?” Understanding Stakeholder Perceptions of LLM-based Chatbots in the Classroom

Abstract.
In recent years, Large Language Models (LLMs) rapidly gained popularity across all parts of society, including education. After initial skepticism and bans, many schools have chosen to embrace this new technology by integrating it into their curricula in the form of virtual tutors and teaching assistants. However, neither the companies developing this technology nor the public institutions involved in its implementation have set up a formal system to collect feedback from the stakeholders impacted by them. In this paper, we argue that understanding the perceptions of those directly affected by LLMS in the classroom, such as students and teachers, as well as those indirectly impacted, like parents and school staff, is essential for ensuring responsible use of AI in this critical domain. Our contributions are two-fold. First, we present results of a literature review focusing on the perceptions of LLM-based chatbots in education. We highlight important gaps in the literature, such as the exclusion of key educational agents (e.g., parents or school administrators) when analyzing the role of stakeholders, and the frequent omission of the learning contexts in which the AI systems are implemented. Thus, we present a taxonomy that organizes existing literature on stakeholder perceptions. Second, we propose the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, which can be used to systematically elicit perceptions and inform whether and how LLM-based chatbots should be designed, developed, and deployed in the classroom.
ChatGPT; Large Language Models; LLM-powered bots; Perceptions; Education
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Social and professional topics Socio-technical systemsccs: Human-centered computing
1. Introduction
Since their emergence, large language models (LLMs) have spread rapidly across many sectors, including education. Using deep learning techniques, LLMs can be trained on extensive text datasets, allowing them to generate high-quality natural language text and demonstrate contextual understanding, which makes them appealing to a wide range of user groups (Watson and Romic, 2024). Among these transformer-based models, the Generative Pre-trained Transformer (GPT) series has emerged as a prominent leader due to its enhanced capabilities, scalability, and ability to operate across multiple channels and use cases, among other advantages (Bird et al., 2023; Garcia-Mendez et al., 2024; Bansal et al., 2024). In November 2023, only one year after the initial public release of ChatGPT, OpenAI announced that 100 million people were using its chatbot each week (Porter, [n. d.]; Lieb and Goel, 2024).
The rapid rise of generative AI, especially LLM-powered chatbots, presents both opportunities and challenges for educational institutions, prompting them to take a stand on their use (Woodruff et al., 2023). In this context, 2023 proved to be a pivotal year for AI in education. In the United States, the year began with school districts rushing to address the sudden emergence of ChatGPT, with some implementing bans that were later reversed (Rosenblatt, 2023). The year ended with the technology being integrated into school curricula, as evidenced by over 40 districts and 28,000 students and teachers testing Khan Academy’s LLM-powered tutor and teaching assistant, Khanmigo (Khan Academy, 2023), in classrooms (Yamkovenko, 2024).
While some have welcomed this technology into the classroom with excitement and openness, occasionally comparing the adoption of ChatGPT to the introduction of calculators and computers (Shoufan, 2023), its use has also sparked concerns and mixed reactions among stakeholders, particularly students and teachers (Engle, 2023; Ravaglia, 2024).
The position we take in this paper is that stakeholders—individuals or groups impacted by the adoption of technology directly or indirectly—should play an integral role in the design, development, implementation, and oversight of AI in socially-salient contexts. Further, we posit that, when it comes to AI systems such as LLM-based chatbots in educational contexts, the number of agents considered as stakeholders should be extended beyond students and teachers to include families, school administrators, ed-tech professionals, and government institutions. We will explore a specific dimension of stakeholder engagement, namely, understanding stakeholder perceptions, which involves eliciting and analyzing what goals the stakeholders pursue and how they view the AI system (e.g., in terms of opportunities and concerns). Understanding perceptions enables key dimensions of the responsible use of technology, including trustworthiness, safety, fairness, transparency, and usability (Department of Education, Office of Educational Technology, 2023).
1.1. Motivating example: LLM-based chatbots in New York City public schools
In March 2023, OpenAI released a new model, GPT-4. As a launch partner, Khan Academy had early access to the model, and on March 14, they released their GPT-4-powered tutoring bot, Khanmigo, in a limited pilot period (Yamkovenko, 2024). Newark Public Schools was one of the first school districts in the U.S. to pilot the tool (Yamkovenko, 2024). The model initially followed a fee-per-student structure, and later became available for free, thanks to a partnership between Khan Academy and Microsoft, which donated its Azure AI-optimized infrastructure. This helped address one of the key obstacles that educational innovations face, namely securing funding (Ravaglia, 2024; Engle, 2023).
Following this initiative, similar efforts were launched in New York City. In September 2023, the NYC Department of Education (DOE) partnered with Microsoft to create a custom AI-powered teaching assistant designed to answer students’ questions and provide personalized, real-time feedback in the classroom (Donaldson, 2023). The tool was first piloted in high school computer science courses, after which, a feature called Microsoft’s Math Solver was incorporated to the tool to improve accuracy on math problems in 15 public schools (Donaldson, 2023). At the same time, teachers in Northern Brooklyn were being trained on how to use YourWai, an AI teaching assistant created by the company Learning Innovation Catalyst, which was later on criticized for displaying made-up reviews on their website (Bardolf and Edelman, 2023). More recently, NYC schools were forced to withdraw a \$1.9 million proposal for an AI-based reading tutor, Amira, designed to assess students’ reading abilities. The proposal was halted by the city Comptroller until the DOE established clearer policies and student privacy safeguards (Zimmerman, 2024). The tool was reportedly used by 46,000 students across 162 NYC public schools (Zimmerman, 2024).
In the article “Would You Want an A.I. Tutor?”, the New York Times invited students to submit comments, sharing their perspectives on using Khanmigo in the classroom (Engle, 2023). [^1] The title of our paper makes a direct reference to Engle’s New York Times article (Engle, 2023), which inspired our work. Opinions were mixed, with some students expressing a range of concerns, such as the lack of human interaction: “I think having A.I. tutoring is not a great idea, because human interaction is a really good way to learn”; the risk of over-reliance: “I’d say there should be a limit that students should know when using A.I. for help. When an assignment is obviously involving critical thinking, you need to exercise your brain by thinking outside the box”; or concerns about accuracy: “In some cases, one may ask a simple question to Artificial Intelligence, and the answer may be correct, but how would we know for certain? Where is the evidence of said ‘correct answers’?”. Positive perceptions were also expressed, with some students appreciating the wide knowledge base of the chatbot: “I think that A.I. can be helpful because the information A.I. have could be endless”; its accessibility: “with a AI tutor you can get help anywhere in seconds”; and the personalized tutoring feature: “I am 100% in the belief that an AI tutor that’s personalized for students and maybe in the future, personal tutoring profiles would be an extreme benefit for students and teachers in terms of saving time and learning more.”
1.2. Research gap: Lack of systematic understanding of stakeholder perceptions of educational chatbots
Systemic concerns regarding the lack of stakeholder engagement have been raised in the literature. For example, Zeide (2019) writes that “[…] school procurement and implementation of these systems are rarely part of public discussion. Students and parents cannot observe, let alone challenge, the many decisions that shape academic outcomes and subsequent life trajectories. Teachers and school officials cannot assess the accuracy, efficacy, and fairness of individual determinations or group outcomes. Private companies, rather than community members and public servants, set pedagogy and policy in practice.” Further, a growing body of research has in recent times emphasized the need for stakeholder-first approaches in the development and deployment of AI systems, as well as the need for actions by governments and civil society to establish more rigorous standards and regulatory measures, grounded in international human rights frameworks, that can hold Big Tech and other powerful actors accountable (Fukuda-Parr and Gibbons, 2021; Bell et al., 2023).
Limited stakeholder involvement contrasts sharply with the recommendations in the U.S. Department of Education’s May 2023 report, “Artificial Intelligence (AI) and the Future of Teaching and Learning: Insights and Recommendations” (Department of Education, Office of Educational Technology, 2023). The report outlines four foundational principles for policy development: Centering People (parents, educators, and students); Advancing Equity; Ensuring Safety, Ethics, and Effectiveness; and Promoting Transparency (Department of Education, Office of Educational Technology, 2023). Its first key recommendation, “human in the loop,” underscores the importance of educators, parents, families, students, policymakers, and system leaders in exercising judgment and control over AI’s role in education.
Ed-tech vendors are also increasingly recognizing the importance of stakeholder engagement. For instance, Microsoft’s Technical Implementation Guide for New York City Public Schools highlights community engagement as a key step for successfully implementing their AI teaching assistant (Microsoft, 2024). The guide emphasizes “Building trust and support with stakeholders” through feedback, input, and addressing concerns.
And yet, while initiatives to implement LLM-powered aids in the classroom have been proliferating, no systematic methodology has been established to assess stakeholder perceptions of the newly introduced technology, either before or after its rollout. Our paper aims to contribute to this literature by illustrating the current landscape of stakeholder perceptions of LLM-based aids in the classroom, and providing a framework for contextualizing these perceptions within the educational setting.
Figure 1. Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE), a framework for assessing and contextualizing stakeholder perceptions of LLM-based chatbots. Co-PACE comprises four components: Stakeholders, Context, Goals, and Perceptions. Starting at the top, one should consider components at each layer before moving to the next layer. The Perceptions component can be used as a checklist to assess the presence of specific perceptions, refining them and adding any that are missing.
1.3. Contributions and paper organization
Initially, we set out to conduct a literature review of stakeholder perceptions of LLM-powered chatbots in the classroom. We were hoping to learn how a school district such as New York City’s might “do things right” — ask the right stakeholders the right questions about whether and how to deploy chatbots in the classroom. However, we quickly realized that there are significant gaps in the literature: (i) only students and teachers have been considered, while other stakeholders (e.g., parents) have been overlooked; and (ii) the context of use of a chatbot in the classroom (e.g., whether it is used for language or math instruction, in elementary school or in a high school) is underspecified. Nonetheless, we were able to distill a taxonomy of perceptions that can be used to characterize stakeholder attitudes towards LLM-based tutors. Our literature review, and the taxonomy that emerged, is the first contribution of this paper. Our second contribution is the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, summarized in Figure 1.
This paper is organized as follows. Section 2 summarizes related work, Section 3 describes our literature review of perceptions of LLM-based chatbots. Section 4 presents the Co-PACE framework, which encompasses the taxonomy of perceptions. Section 5 concludes the work, acknowledges limitations, and outlines future directions.
2. Related Work
2.1. Reviews of perceptions of LLM-based chatbots
An increasing number of studies explores perceptions of LLM-based chatbots in educational contexts, often using sentiment, emotion, thematic, or content analysis.
Leiter et al. (2024) examined ChatGPT’s adoption two months post-launch, analyzing 300,000 tweets and 150+ scientific papers. Their findings revealed a balanced divide in education-related papers, viewing ChatGPT as both an opportunity (e.g., enhancing writing skills and efficiency) and a threat (e.g., to academic integrity).
Mogavi et al. (2024) conducted a qualitative study on early users’ perceptions of ChatGPT in education, analyzing content from four major social media platforms. Their findings indicate widespread use in higher education, K-12, and practical skills training, with productivity, efficiency, and ethics being common discussion topics. While some users saw ChatGPT as a transformative tool for enhancing self-efficacy and motivation, others raised concerns about overreliance on AI, superficial learning habits, and diminished social and critical thinking skills.
Dube et al. (2024) examined university students’ perceptions of the use of ChatGPT in academic activities. Positive perceptions included learning support, concept simplification, study material organization, efficient information retrieval, reliable tutoring, research aid, enhanced motivation, editing convenience, task optimization, problem-solving, data analysis, and mentorship. Negative perceptions highlighted concerns about reduced creativity, privacy and security risks, diminished trust in academic work, unreliable references, and issues with cheating and plagiarism.
Lo (2023) examined ChatGPT’s impact on education, focusing on performance across subjects, functions, and concerns. Performance varied from outstanding to unsatisfactory across domains. For teaching, ChatGPT was primarily used for lesson preparation and assessment, while for learning, it served as a virtual tutor, supporting learning through Q&A and facilitating collaboration. Key concerns included accuracy, reliability, and plagiarism prevention.
Similar to our paper, these works examine LLM-powered chatbots in education. However, some focus exclusively on widely studied stakeholder groups, such as university students (Dube et al., 2024), while those that include less-studied stakeholders (e.g., parents) (Leiter et al., 2024; Mogavi et al., 2024) fail to emphasize recurring perceptions within each group. Furthermore, they report on these stakeholders only during the early adoption phase, specifically within the first few months following ChatGPT’s launch in late 2022. Lastly, some studies (Lo, 2023) analyze LLM adoption solely in terms of its functions, focusing only on negative perceptions or concerns while overlooking the opportunities this technology offers.
2.2. The role of stakeholders in the design and implementation of AI-based systems in education
Several studies have attempted to involve stakeholders in the development and implementation of AI-powered classroom tools. For instance, Chaudhry et al. (2022) mapped transparency requirements for AI-based ed-tech tools. By engaging educators, parents, ed-tech experts, and AI practitioners, they demonstrated that transparency depends on the target audience and evolves across stakeholder groups in educational contexts. They also co-designed a Transparency Index framework, applying and iteratively improving it in a real-world setting where AI-powered educational tools were developed for a training organization.
Luckin et al. (2022) proposed a 7-step AI readiness framework for the education and training sector. They argue that individuals and organizations achieve AI readiness when they can identify: (a) organizational challenges best addressed by data and AI, (b) the context for effective AI application, (c) the goals of AI-based tools, and (d) the stakeholders involved. The authors emphasize that AI readiness requires an active, participatory training process that empowers people to leverage AI to meet their needs.
Finally, Garcia-Mendez et al. (2024) conducted a literature review of LLM-based solutions for generating and evaluating educational materials involving students or teachers in their design or experimental plans. Their goal was to promote collaboration between academia (i.e., researchers and developers) and end-users (i.e., students and educators). The review found that the most common functions performed by LLMs as virtual assistants in these solutions were question generation, answer grading, and code correction and explanation.
Building on the works described above, we advocate for the inclusion and systematic collection of stakeholder perspectives to inform the design and deployment of LLM-powered chatbots in the classroom.
3. A Review of perceptions of LLM-based chatbots
We aimed to analyze stakeholder perceptions of LLM-powered chatbots in educational settings, with a focus on identifying the most recurring perceptions for each stakeholder group. We were guided by two research questions:
RQ1: Who are the key stakeholders and what are their perceptions, as identified in the literature?
RQ2: What are the main gaps in the literature regarding stakeholders and their perceptions?
Table 1. Number of papers per type of study design
Study type # studies
Non-experimental Interviews or open-ended questions surveys 13
Mixed-methods 12
Questionnaires 3
Quasi-experimental Pre-test/Post-test 2
Mixed-methods 1
Pre-Experimental One-shot case study 1
Experimental Mixed-methods 3
Randomized controlled trial 2
Multi-arm randomized trial 1
3.1. Methods
3.1.1. Selection procedure
We used Google Scholar to search for scientific literature on the topics of interest. To be considered, articles had to include the search strings (“Large language models” OR “LLMs” OR “LLM-based tutor” OR “LLM-powered chatbot” OR “ChatGPT”) AND (“perceptions” OR “perspectives” OR “attitudes”) in their title, abstract, keywords, or introduction. For the selection procedure, we limited the number of results to the first 100 articles, using Google Scholar’s relevance sort order. Articles were included if they met the following criteria: (1) discussed LLM-powered bots in education; (2) addressed perceptions of education stakeholders, such as students, teachers, parents, school staff, or education-related professionals; (3) were based on experimental, quasi-experimental, pre-experimental, or non-experimental research (excluding commentaries, secondary data analyses, and literature reviews); and (4) were written in English. Finally, the study must involve LLM-powered chatbots. Studies that only mentioned “AI tool” without specifying whether they referred to large language model-based technology or rule-based systems were excluded. This resulted in the selection of $N=38$ full-text articles for analysis.
3.1.2. Content analysis
After an initial reading of the articles, we identified recurring patterns or themes and developed a codebook to capture them. The codebook was developed combining codes taken from articles that employed existing theoretical frameworks (deductive) and codes that emerged from the articles without available codebooks or that did not develop one (inductive). In the latter cases, we annotated the results and discussion sections of the articles. As for the level of implication allowed, we included both explicit references to concepts and those implied by words or phrases. We allowed for flexibility by adding codes throughout the iterative readings.
Hence, we coded for the frequency, that is, we examined the presence of a selected code in the articles. We identified whether each code or concept reflected negative perceptions or concerns, positive perceptions or opportunities, or neutral perceptions. Additionally, we identified the specific stakeholder group(s) from which each code or concept originated. The validity of the coding process was ensured through multiple rounds of coding, allowing for refinement and consistency across the analysis.
3.2. Findings
3.2.1. Basic summary information
We now summarize the basic break-down of the $N=38$ articles included in our analysis. As shown in Figure 2(a), the majority of the included articles had samples from Asia ($N=21$), followed by Europe ($N=20$), North America ($N=18$), Oceania ($N=5$), Latin America ($N=4$), and Africa ($N=4$). (Note that several articles included samples from multiple continents.) When the location of the participants was not specified, the location of the authors’ university was selected as a proxy.
(a) continents
(b) educational contexts
Figure 2. Break-down of statistics of the analyzed articles ($N=38$).
Regarding the contexts of application of the LLM-based chatbots, Figure 2(b) illustrates that the most frequent educational context of the studies was STEM ($N=19$, 50%), followed by English-as-foreign-language (EFL) ($N=5$, 13%) and Humanities ($N=1$). A considerable number of studies did not specify the educational context ($N=13$, 34%).
Table 1 presents a break-down by type of study design. Most of the articles employed non-experimental study designs. 13 studies used interviews or surveys with exclusively open-ended questions, most of which were semi-structured. 12 studies used mixed methods, including surveys with both open- and closed-ended questions, as well as less common methodologies, such as having participants draw a visual representation of their perception of ChatGPT followed by a survey (Ding et al., 2023). Three studies used quantitative questionnaires. Three articles used quasi-experimental research designs, two of which used pre- and post-test methodology, and one, mixed methods. One study was pre-experimental, using a one-shot case study design. Among the experimental designs, the most common was mixed methods, followed by randomized controlled trials (RCT) and multi-arm randomized trials.
3.2.2. RQ1: Key stakeholders and their perceptions
Figure 3. Stakeholders whose perceptions were analyzed in the articles ($N=38$).
As shown in Figure 3, students were the most studied stakeholder group, appearing in 31 of the 38 (82%) studies, 27 of which involved University-level students and 4 involved K-12 students. Teacher perceptions were analyzed in 12 (32%) studies. Additionally, perceptions of industry professionals were considered in 1 study, albeit without explicitly specifying whether these were ed-tech professionals or learners. No studies included any other stakeholders, such as parents or school administrators.
We identified different types of perception—positive (opportunities/benefits), negative (concerns), and neutral— in the articles we analyzed. We organized these perceptions into five categories: Attributes, Impact on Processes or Outcomes, Functions, User Experience, and Ethical and Societal Implications. Figure 4 summarizes results of our analysis, showing the total number of mentions of perceptions in each category separately for two stakeholder groups, Students (Figure 4(a)) and Teachers (Figure 4(b)). Complete results, with frequencies for each perception type within each category, are provided in Appendix A. We present additional details about each perception category in the remainder of this section. Note that, while we do not provide quantitative details about industry professionals as a stakeholder group in either Figure 4 or Appendix A, because they were only covered in one article, we do include them when discussing qualitative findings.
Overall, the perceptions of both students and teachers were more frequently positive than either negative or neutral: with 58% of all perceptions were positive and 33% were negative for students vs. 51% positive and 38% negative for teachers. Additionally, while both students’ and teachers’ perceptions focused heavily on the ethics of LLM-based chatbots, particularly in a negative or concerned light, students also expressed views on UX, whereas teachers concentrated on the functions of LLM-powered bots. In contrast to the ethics of LLM-based bots, the topics of UX and functions were primarily addressed by their respective stakeholders in terms of opportunities or positive aspects.
Attributes.
This category refers to perceptions of the characteristics of LLM-based chatbots. The papers we surveyed discussed the perceptions of students, teachers, and industry professionals regarding these attributes. The key attributes identified included: Convenience (also referred to as “accessibility”, “responsiveness” or “speed”), Accuracy (also referred to as “confidence and trust in LLMs”), Knowledge base, Anthropomorphism (also referred to as “human-like qualities”, “perceived humanness” and “Natural Language Processing (NLP) capabilities”), and various other, less popular technical specifications, such as character limit and code editing. Perceptions of these attributes were mixed, with each being viewed both positively and negatively, except for Convenience, which was consistently seen as an advantage. Additionally, no neutral perceptions were found regarding these attributes.
The most frequent positive perception among students and teachers is Convenience. For example, students in Kazemitabaar et al. (2024) appreciated CodeAid’s “24/7 availability”, whereas teachers in Mohamed (2024) highlighted the bot’s “real-time interaction and feedback”.
In contrast, the most recurring concern for students, teachers, and industry professionals is Accuracy. For example, students in Weber et al. (2024) were concerned with the accuracy of answers provided by LLMs, while teachers in Ghimire et al. (2024) stated that generative AI can produce “incorrect or fabricated results.” However, an industry professional in Chen et al. (2024) expressed that although she had “low trust in ChatGPT”, she felt more confident after using it, as it helped her understand what she needed to figure out and made her in general faster.
Impact on Processes and Outcomes.
This category refers to perceptions related to teaching and learning processes and their results. These perceptions were most commonly expressed by students and teachers, and include: Efficiency, Student motivation, Performance expectancy (a concept from UTAUT (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), see Appendix B for a summary of these theories), referring to improved student learning and outcomes, Cost-effectiveness, as well as Critical thinking and problem solving skills. Among these, Performance expectancy, Cost-effectiveness, and Critical thinking and problem solving skills generated mixed perceptions. No neutral perceptions were found regarding these impacts.
Efficiency was the most commonly cited positive perception among students, teachers, and industry professionals. For example, students in Kazemitabaar et al. (2024) stated that AI “helps them work more efficiently,” while teachers in Jansson et al. (2024) reported that “not having to calculate the problem themselves may free up time and make them more efficient as coaches, enabling them to help more students.” Similarly, industry professionals in Chen et al. (2024) noted that “with LLMs, human time and energy could be liberated for more high-level tasks.”
The most common negative perception across students and teachers was the impact on Critical thinking and problem solving skills. For example, a student in Woithe and Filipec (2023) noted that while ChatGPT “is a tool that’s very useful when learning,” they were “still a bit scared [that] I’ll become incompetent to think on my own.” Teachers in Ghimire et al. (2024) also expressed concern, stating that LLMs “decrease critical thinking.”
(a) students
(b) teachers
Figure 4. Summary of perceptions of two major stakeholders, students (4(a)) and teachers (4(b)), reported in the articles in our literature review. Here, Impact refers to “Impacts on Processes and Outcomes,” UX refers to “User Experience,” and Ethics refers to “Ethical and Societal Implications.” See Appendix A for a detailed break-down.
Functions.
This category refers to perceptions related to uses of LLM-powered technology in the classroom, including Assessment, Personalized tutoring, Material creation, Writing assistance, Brainstorming assistance, Opportunity for practice, Tool for complementing traditional teaching, Facilitating conditions (a construct in UTAUT that describes the degree to which an individual believes an organization’s and technical infrastructure support the use of the system (Venkatesh et al., 2003)), as well as Other functions. Some of these perceptions were viewed both positively, negatively, and neutrally, except for Tool for complementing traditional teaching, which was only perceived neutrally.
The most recurring positive perception about functions across all stakeholders was Personalized tutoring. Students recognized “the potential for personalized learning support” (Chan and Hu, 2023) and reported that “ChatGPT excels in breaking down intricate concepts into more understandable terms” (Zimotti et al., 2024). Teachers also emphasized the value of “Personalized teaching/24 hours Teaching Assistant (TA) access” (Ghimire et al., 2024), while industry professionals underscored that LLMs “support human effort” (Chen et al., 2024).
The most common negative perception regarding functions among students and teachers related to Facilitating conditions. For students, Ogbo-Gebhardt and Ogbo (2024) highlighted a lack of institutional support on GenAI usage (Chan and Hu, 2023; Ogbo-Gebhardt and Ogbo, 2024). For teachers, concerns included banning LLM-based tools (Lau and Guo, 2023) and the “lack of teacher support for using ChatGPT as an educational tool” (Iqbal et al., 2022). Similarly, the most commonly cited neutral perspective among students involved Facilitating conditions. For example, students in Woithe and Filipec (2023) stated, “[…] addressing these issues, not by banning ChatGPT, but by implementing regulatory measures and ethical guidelines for its use.” Teachers, on the other hand, emphasized the need for support in implementing LLM-based bots in the classroom. As one teacher in Iqbal et al. (2022) noted, “I feel like there should be more training available on how to use ChatGPT in the classroom. It would really help if we could get more help from the administration in this.”
User Experience (UX)
This category of perceptions relates to how individuals feel about interacting with LLM-powered bots. It includes concepts such as perceived Usefulness, Ease of use, Intention to use, Hedonic motivation, Well-being, Perceived importance, and Human interaction loss. Many of these—like Usefulness, Ease of use, Intention to use, and Hedonic motivation—originated in the Technology Acceptance Model (TAM) (Davis, 1989) and evolved in the Unified Theory of Acceptance and Use of Technology (UTAUT) (Venkatesh et al., 2003) and UTAUT2 (Acosta-Enriquez et al., 2024), often with different names (see Appendices B and C). For instance, perceived Usefulness in TAM is called Performance expectancy in UTAUT (Sair and Danish, 2018), but we retained the term perceived Usefulness due to its use in studies beyond these frameworks. Similarly, TAM’s Ease of use evolved into UTAUT’s Effort expectancy, reflecting the role of technical skills in technology acceptance (Venkatesh et al., 2003). Except for Human interaction loss, which is perceived only negatively, and Hedonic motivation, viewed positively, negatively, and neutrally, all other concepts are perceived both positively and negatively.
The most commonly reported positive perception among students was perceived Usefulness. Findings in Elkhodr et al. (2023) show “a positive perception of ChatGPT as useful”, while Lee et al. (2023) state that “Learners’ perception of the AI Teaching Assistant (TA) is on par with that of human TAs in terms of […] helpfulness.” The most positive perception among teachers was Ease of use. For example, Jansson et al. (2024) reports that “The coaches expressed surprise about how friendly and supportive the AI bot was,” while teachers in Iqbal et al. (2022) mentioned that ChatGPT is “extremely user-friendly.” Industry professionals reported positive perceptions of both Usefulness and Intention to use (Chen et al., 2024).
The most common concern among both students and teachers is Human interaction loss. For example, a student in Ogbo-Gebhardt and Ogbo (2024) disclosed that “she felt a bit of anxiety considering how increased adoption of LLM tools could lead to increased social isolation,” while Chan and Hu (2023) reported that “in academic institutions and education, some were concerned that the widespread use of AI might affect the student–teacher relationship, as students may be ’disappointed and lose respect for teachers.”’ Teachers also expressed concerns about the “lack of human connection” (Mohamed, 2024). Interestingly, Ease of use was equally as concerning as Human interaction loss among teachers, with some mentioning that “they found it difficult to use” (Iqbal et al., 2022) when discussing ChatGPT.
The only neutral UX perception in the literature is Hedonic motivation, a UTAUT2 concept referring to the fun or pleasure of using a technology. For instance, Lieb and Goel (2024) observed that “the neutral hedonic evaluation indicates that users had more neutral feelings about how interesting or exciting it was to use NewtBot.”
Ethical and Societal Implications.
This category covers the anticipated impact of introducing LLM-based bots in schools, including Social justice, Job replacement, Reliance, Ethical and privacy concerns, Academic integrity, Social influence (a construct from UTAUT that captures the impact of family and friends on technology acceptance), Workplace changes, and Learning changes. While all were viewed positively and negatively, the last four also had neutral perceptions.
The most frequently mentioned positive perception among students is Reliance on LLM-based bots. For instance, Smith et al. (2024) noted that “few students described behaviors that appeared to be intentional cheating (i.e.,’generating complete answers’),” demonstrating a healthy level of reliance on LLMs. Similarly, Kazemitabaar et al. (2024) reported that “although some students, like S14, felt that they ’over-relied on it too much rather than thinking,’ many students displayed signs of self-regulation.” This positive perception is also evident among industry professionals. A participant in Chen et al. (2024) compared ChatGPT (GPT-4) and NetLogo, commenting that ChatGPT “assumed what I wanted it to do, whereas this one makes you specify your assumptions.” They preferred NetLogo Chat’s approach because “it makes you think about the code more.” For teachers, the most common positive perception was social justice, encompassing “equity,” “accessibility,” and “cultural understanding.” Teachers in Mohamed (2024) noted that ChatGPT could become a “tool for enhancing language exposure and cultural understanding […] and a tool for enhancing language accessibility and equity.” Similarly, teachers in Lau and Guo (2023) stated that “AI may improve equity and access.”
The most frequent concern among students was also related to Reliance, which, in this case, takes the shape of overreliance. This concern is reflected in statements such as, “I feel that everyone might gradually lose their ability to think, as people won’t be willing to think anymore. They’ll just rely on machines to do the thinking for them” (Zhang et al., 2024), or “While students are specifically informed that AI can make mistakes and the exam keys are even provided, almost half of the students still agree with all the answers provided by ChatGPT, regardless of whether the answers are correct or incorrect” (Ding et al., 2023). Similar concerns about Reliance were also shared by industry professionals. The most common concern among teachers is Academic integrity, often framed as “plagiarism” or “AI-assisted cheating.” For example, teachers in (Ghimire et al., 2024) expressed concern about “focus on product over process,” while Iqbal et al. report that most teachers surveyed “found ChatGPT would be used by students for cheating, and it will make them lazy.”
Finally, the most common neutral perceptions regarding the Ethical and Societal Implications of LLM-powered tools were related to Changes in the workplace for students and both Changes in the workplace and Changes in learning for teachers. Regarding the former, Rogers et al. found that “Almost all students agreed that ChatGPT will be integral to their careers,” while Bernabei et al. reported that “students recognized the practical relevance of using such tools in a university context, as they believed it would better prepare them for the future working world, where AI is anticipated to play a prominent role.” As for the latter, a teacher in (Mohamed, 2024) stated, “I see ChatGPT playing an increasingly central role in EFL education in the future.”
3.2.3. RQ2: Gaps in the literature
Stakeholder representation.
Our literature review highlighted a lack of representation of stakeholders beyond students and teachers, such as parents and broader school staff. Figure 3 shows that students appeared as the most recurring stakeholders, followed by teachers. This finding mirrors a conclusion from a literature review on robot tutors by Smakman and Konijn (2020), who also noted the lack of parental representation. It further underscores the need identified by Louie et al. (2022) to include “key stakeholders (students, parents, and teachers) as essential to ensure a culturally responsive robot.” Finally, it echoes the concerns raised by Zeide (2019) that “school procurement and implementation of these systems are rarely part of public discussion.”
Context representation.
The second gap highlighted by our literature review was the inadequate representation of the educational context, particularly in terms of discipline and educational levels of the students’ sample. Figure 2(b) illustrates the contexts of application of LLM-based chatbots, with the most frequent ones being STEM, followed by studies with unspecified contexts. Additionally, a lack of diversity in educational levels could be observed in Figure 3, where 27 out of the 31 studies involving students focusing on university-level students and only four studies involving K-12 students. This is a critical issue because, while university students can independently choose whether to use AI technologies, K-12 students typically do not have a say in the implementation of these technology within their curricula (Singer, 2023; of Palm Beach County, 2024). This underscores the need for further research on perceptions across a wider range of educational contexts.
Homogeneity in assessment of perceptions.
Finally, the wide variety of theories, scales, and terms used across the literature to assess and name perceptions highlights the need for standardization in these aspects (see Appendices B and C). For example, in Ogbo-Gebhardt and Ogbo (2024), the theme of trust in and accuracy of the responses generated by LLM-based bots is considered a Facilitating condition (UTAUT), whereas in multiple other papers (Kazemitabaar et al., 2024; Ding et al., 2023; Yilmaz and Yilmaz, 2023) it is treated as a separate concept. This lack of unified terminology makes it difficult to develop a clear understanding of how stakeholders perceive LLM-based technology. We advocate for the standardization of terms in the field.
4. Proposed Framework
We propose a methodological framework, Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) (Figure 1), to guide policymakers, researchers, and developers in adopting AI-based chatbots more responsibly in educational contexts by incorporating stakeholder perceptions. Co-PACE addresses systematic gaps in the literature while supporting the iterative refinement of its nested taxonomy of perceptions.
The creation of Co-PACE is articulated around the perceptions of the different actors involved in the use of AI chatbots in education. Specifically, it is based on a stakeholder-first approach that aims to support key agents in designing and implementing transparent, compliant, and accountable AI systems (Bell et al., 2023). The second component is the goals that stakeholders have in mind when designing and/or implementing AI chatbots in educational settings. The context component was also added to address the lack of specification of the educational settings in which LLM-based chatbots are used, as shown in the literature review. And finally, the taxonomy of perceptions was added as a ’checklist’ for policymakers, technologists, and researchers to consult when assessing the presence of specific perceptions in a given setting and is designed to be refined over iterations through future studies. In the remainder of this section, we will describe each component in depth: stakeholders, context, goals, and taxonomy of perceptions, and explain how the framework should be used.
4.1. Stakeholders
We highlighted how the introduction of LLM-based chatbots, such as Khanmigo, in the classroom could be seen as a top-down decision-making process, where partnerships between private companies and governmental institutions or school districts often make decisions on behalf of direct stakeholders, such as teachers and students, or indirect stakeholders, such as parents and school staff. This echoes a concern already highlighted in the literature regarding how private companies, more often than community members and public servants, set pedagogy and policy in practice (Zeide, 2019). In contrast to this tendency, our framework prioritizes educational stakeholders: we identified students and teachers as the most recurring stakeholders in the literature; however, the systematic gaps we found underscore the need to assess the perceptions of other key stakeholders, such as parents, school staff, ed-tech professionals, and government agencies.
4.2. Context
As pointed out in Section 3.2.3, most studies in the literature have been conducted in the field of higher education. This underscores the need to assess the perceptions of stakeholders at other educational levels, such as younger students who lack the autonomy to make independent decisions, yet are still expected to use LLM-based chatbots as part of their curricula. The gaps also indicate that most studies focus on the STEM field, with many others having unspecified areas of application. We advocate for the need to assess perceptions of LLM-based chatbots across various disciplines.
4.3. Goals
The goals of soliciting and incorporating stakeholders’ perceptions when developing and deploying AI systems are diverse. In our framework, we propose goals drawn from the literature on responsible design of AI systems and organize them into two focus areas: those focused on system behavior and those focused on educational practice. When the focus is on system behavior, the goals are related to designing these tools or assessing whether they operate according to responsible AI key principles. We align with the existing literature on transparency, which asserts that the goal of a transparent design must begin with identifying the stakeholders (Bell et al., 2023; Chaudhry et al., 2022). Bell et al. (2023) established six categories to assist technologists in designing transparent, regulatory-compliant AI systems: validity, trust, learning and support, recourse, fairness, and privacy. Chaudhry et al. (2022) additionally emphasizes safety as a crucial component of transparency, especially in education, where mishaps in AI-powered technologies can often go unnoticed unless reported to the relevant authorities by a teacher or a student. Finally, although Louie et al. (2022) discuss this in relation to social robots rather than LLM-based chatbots, they distance themselves from the traditional designer-determined approach, advocating for the inclusion of culturally and linguistically diverse stakeholders as co-designers and expert informants in the design process. They promote participatory design, reporting that such inclusion leads to more engaging designs, improved user experience, and greater academic gains and ownership of technology-based learning.
In terms of pedagogical goals, the focus is on embedding the responsible principles in educational practices and relating them to specific learning outcomes (Domínguez Figaredo, 2020; Fu and Weng, 2024). For example, the in-depth analysis of Jurenka et al. (2024) refers to the specific learning practices embedded in the LearnLM tutor that the authors evaluated to improve the pedagogical capabilities of the AI system while promoting more responsible and effective learning experiences, including: evaluative practices, feedback on procedural problems, grounding of learning materials, adaptivity to learner level, engagement, etc. In Co-PACE, we have presented some of the goals that could be associated with implementing responsible practices with AI systems in education, and this is an area that practitioners could expand according to their vision and needs.
4.4. Taxonomy of perceptions
Our survey identified a gap in the inconsistent terminology used for similar concepts. To address this, our proposed taxonomy serves as a checklist for researchers, ed-tech professionals, and policymakers to assess, refine, and expand stakeholder perceptions, promoting greater consistency in perception research.
4.5. Putting the Co-PACE framework into practical use
The Co-PACE framework is designed to assist policy makers, researchers and technologists in two areas:
  1. (1)
    Elicit stakeholders’ perceptions to create new or refine existing recommendations to ensure that AI systems are human-centered and responsibly adopted.
  2. (2)
    Guide researchers and practitioners to address gaps in the literature on LLM-based chatbot perceptions through future studies.
In practice, the application of Co-PACE follows the sequence shown in Figure 1: starting with the first component, “stakeholders,” one should consider one or more of the bubbles before moving on to the next component. For example, policymakers and technologists of an ed-tech company might want to assess the perceptions of the parents of elementary school students of LLM-based chatbots when they are applied to humanistic subjects with the goal of improving or assessing their perceived effectiveness. Upon reaching the final component, the taxonomy, one could use it as a checklist to identify the presence of specific perceptions. In this example, when discussing effectiveness, perceptions related to Impact on Processes or Outcomes are likely to arise. Another example involves assessing teachers’ perceptions of LLM-based chatbots applied to English-as-a-Foreign-Language. The goals might include improving their experience in using this technology, as well as managing potential risks. When going through the taxonomy “checklist,” likely perceptions related to User Experience (UX) and Ethical and Societal Implications, such as cultural understanding (included in the “Social justice” perception), might arise. Co-PACE is an open framework, and so any missing elements in any component that emerge from conducting studies in less explored scenarios could be added.
5. Conclusion, Limitations, and Future Work
LLM-based chatbots are being piloted in schools, often through partnerships between private companies and school districts. However, feedback from direct users (students and teachers) and indirect users (parents and school staff) has been largely absent before and during rollout, aside from a few informal efforts. To address this gap, we conducted a literature review on stakeholders’ perceptions of LLM-based chatbots in education, identifying key positive, negative, and neutral views. Students generally perceive these chatbots as useful and efficient but raise concerns about accuracy. Teachers value their role as personalized tutors but are primarily concerned with maintaining academic integrity. Neutral perceptions from both groups describe chatbots as complementary to traditional teaching. The review also revealed gaps in the literature: (i) underrepresentation of key stakeholders, such as parents; (ii) limited exploration of diverse educational contexts, including various levels and disciplines; and (iii) inconsistent assessment of perceptions, with varying theories, scales, and terminology applied to similar concepts. We then proposed the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, incorporating a taxonomy of perceptions from the literature. Co-PACE is designed to help policymakers and technologists assess stakeholder perceptions throughout the development and deployment of AI systems in education, promoting more responsible and trustworthy processes.
Limitations
This article has limitations that must be considered when interpreting the findings. First, the organization of stakeholders’ perceptions into the taxonomy, as well as the assignment of values (positive, negative, or neutral) to them, was based on the researchers’ judgment during the content analysis process. Future iterations could help make the taxonomy more rigorous. Second, the literature review was limited to articles published on Google Scholar. Including articles from databases such as Scopus, ISI Web of Science, and other cross-references could add weight and rigor to the findings related to stakeholders’ perceptions of LLM-based chatbots in education. Lastly, the in-depth analysis of the literature was based on empirical studies conducted globally. Narrowing the scope to a particular region or tool could provide deeper insights, as perceptions may be influenced by cultural and social norms, socio-economic status, and other factors related to the context and background of the stakeholders.
Future work
There are several promising directions for future research. First, the stakeholder-first approach described here lays the foundation for future studies investigating neglected stakeholders’ perceptions across a range of overlooked contexts. Second, a repository of findings derived from the application of this framework could be created, to act as a reference for technologists and policymakers in future technology design processes and implementations.
Appendix A Supplementary statistics
Complete break-down of perceptions, for all perception types in our taxonomy, for two main stakeholder groups: students and teachers. These results complement the summary in Figure 4 in the main body of the paper.
Appendix B Models of Perception
Complete breakdown of models and theories conceptualizing attitudes, beliefs, and/or behaviors toward technology, as they appear in the literature, to investigate stakeholders’ perceptions of LLM-based chatbots. The breakdown supports the thesis outlined in the third gap identified in the literature in Section 3.2.3, namely, the lack of homogeneity in the assessment of perceptions and the need for standardization in how perceptions are conceptualized.
Technology Acceptance Model (TAM) The Technology Acceptance Model (TAM) was first developed in 1989 by Davis and Davis et al.) as a tool to predict the likelihood of adopting a new technology within a group or an organization. Originating in the psychological theory of reasoned action and theory of planned behavior (TPB) (Kowalska-Pyzalska, 2023), TAM suggests that an individual’s beliefs, attitudes, and intentions can explain their adoption and use of technology. TAM evaluates the influence of four internal variables on actual technology use, namely, Perceived Ease of Use (PEU); Perceived Usefulness (PU); Attitude Toward Use (ATU), and Behavioral Intention to Use (BI) (Davis, 1989; Turner et al., 2010; Kowalska-Pyzalska, 2023).
Unified Theory of Acceptance and Use of Technology (UTAUT) The Unified Theory of Acceptance and Use of Technology (UTAUT) originates as an extention of TAM and was originally used to explain user behavior and intentions in information system contexts (Venkatesh et al., 2012; Kowalska-Pyzalska, 2023; Shakib Kotamjani et al., 2023). The UTAUT model comprises several key constructs that affect users’ acceptance and use of technology. These are: Performance Expectancy, Effort Expectancy, Social Influence and Facilitating Conditions. According to Venkatesh et al., the Effort Expectancy factor derives from the Ease of Use factor included in the Technology Acceptance Model (TAM). Additionally, the UTAUT model includes four other moderating variables, i.e., Gender, Age, Experience, and Voluntarism of Use, that may influence the four main constructs (Venkatesh et al., 2012; Shakib Kotamjani et al., 2023).
Extended Unified Theory of Acceptance and Use of Technology (UTAUT2) The Extended Unified Theory of Acceptance and Use of Technology (UTAUT2) is an extension of the prevision one, as it incorporates additional constructs such as hedonic motivation, price value, habit, and usage intention, besides the original ones (Acosta-Enriquez et al., 2024). Additionally, in UTAUT2, voluntarism of use was dropped as moderator. The predictive ability of UTAUT2 theory is found to be much higher in comparison to UTAUT (Tamilmani et al., 2021).
ABC Model of Attitudes The ABC Model of Attitudes (Eagly and Chaiken, 1998) conceptualizes attitudes as being composed of three elements: Affect, Behavior, and Cognition. Affect refers to the individual’s feelings or emotional response toward an object. Behavior reflects the individual’s intentions or actions toward an object or situation. Cognition refers to the beliefs or thoughts an individual holds about an object (Ajlouni et al., 2023). These three components collectively constitute an individual’s attitude toward an object, person, issue, or situation.
Cultural-Historical Activity Theory (CHAT) The Cultural-Historical Activity Theory (CHAT) offers a perspective for examining the intricate and dynamic systems in which AI is integrated into education. When CHAT is applied to technology or AI in education, it doesn’t only consider the student-technology relationship, but the whole activity system (subject, object, tools, rules, community, and division of labour) within the wider socio-cultural and historical context (Carbonel and Jullien, 2024). The underlying concept is that when a new element, such as a new technology, is introduced into the activity system, it often leads to tensions and contradictions. The model can help examine the system and begin envisioning new approaches.
Situated Expectancy-Value Theory (SEVT) Renamed from the original expectancy-value theory, the Situated Expectancy-Value Theory (SEVT) highlights that students’ expectancy-value beliefs—how well they believe they will perform on an upcoming task—are situation-specific, meaning they vary across contexts and are influenced by situational characteristics (Eccles and Wigfield, 2020). SEVT is built on two main components, subjective task value and expectation of success. Subjective task value consists of four parts that influence achievement-related choices and performance:
  • Attainment value: how important the task is to the individual.
  • Intrinsic value: how interesting or enjoyable the task is.
  • Utility value: how useful the task is to the individual.
  • Relative cost: the perceived cost of engaging in the task.
  • Expectation of success refers to how likely the individual believes they are to succeed at the task (Eccles and Wigfield, 2020).
Extended three-tier Technology Use Model (3-TUM) The extended Three-tier Technology Use Model (3-TUM) is based on the original 3-TUM, which integrates multidisciplinary perspectives—including motivation, social cognitive theory (SCT), theory of planned behavior (TPB), and the technology acceptance model (TAM)—to investigate attitudes toward information technology (Liaw, 2007). Individual attitudes are divided into three tiers: the individual characteristics and system quality tier, the affective and cognitive tier, and the behavioral intention tier. The individual characteristics tier includes factors such as self-efficacy, hedonic motivation, and self-regulation, while the system quality tier encompasses environmental factors. The affective and cognitive tier includes perceived satisfaction, perceived usefulness, and performance expectancy (Cai et al., 2023; Liaw, 2008). The constructs of performance expectancy and hedonic motivation are inspired by the Extended Unified Theory of Acceptance and Use of Technology (UTAUT2). The extended model adds an additional tier: learning effectiveness (Cai et al., 2023).
Mitcham’s philosophical framework of technology Mitcham’s (Mitcham, 1994) philosophical framework of technology presents a comprehensive view of technology, emphasizing that technological knowledge and volition, rooted in human nature, lead to technological activities and the creation of concrete technological objects. The framework accommodates the full historical and conceptual range of technology—from ancient to modern and from simple to complex forms—without restriction to any specific type. Using this framework, the Mitcham Score was developed to classify students’ descriptions of technology through four key elements, namely Objects, Activities, Knowledge and Volition (Svenningsson, 2020).
Specific Forms of Digital Competence Needed to use ChatGPT The Specific Forms of Digital Competence Needed to use ChatGPT (Xiao and Zhi, 2023) is a conceptual model based on (Instefjord and Munthe, 2017)’s work on the integration of professional digital competence in initial teacher education programmes. The framework comprises three main components: Technological proficiency (e.g., Be aware of the features of ChatGPT, Understand how ChatGPT works), Pedagogical compatibility (e.g., Think about and plan ways to use ChatGPT to enhance or transform language teaching and learning tasks, Implement tasks that use ChatGPT) and Social awareness (e.g., Have a critical awareness of the drawbacks of ChatGPT and consider them when planning and implementing tasks, Inform learners of the risks, ethical issues, and drawbacks of ChatGPT).
Appendix C Perception Scales
Breakdown of scales assessing attitudes, perceptions, feelings and/or behaviors toward technology as they appear in the reviewed studies. Similar to the breakdown of models, the present lists supports the thesis outlined in the third gap identified in the literature: the lack of homogeneity in the assessment of perceptions and the need for standardization in how perceptions are evaluated.
In addition to theoretical frameworks, several studies have assessed perceptions of LLM-based tutors using existing scales, either fully or partially:
Relevance of Science Education (ROSE) The Relevance of Science Education (ROSE) questionnaire (Schreiner and Sjøberg, 2004) is a research instrument created as part of an international comparative project meant to shed light on students’ experiences, interests, priorities, images and perceptions that are relevant to the learning of science and technology and their attitudes towards these subjects. The questionnaire contains 247 questions across 10 sections and assesses students’ interest in, attitude towards, and experiences in science and technology, as well as their opinion about environmental challenges and career aspirations.
Short Version of the User Experience Questionnaire (UEQ-S) The User Experience Questionnaire (UEQ) aims at collecting end-users’ feelings, impressions, and attitudes that arise when experiencing a product. It consists of 26 items grouped into 6 scales, namely Attractiveness, Perspicuity, Efficiency, Dependability, Stimulation and Novelty (schrepp2017design). Schrepp, Hinderks, and Thomaschewski later developed a Short Version of the UEQ, reducing it to 8 items grouped into two scales, which measure two meta-dimensions: pragmatic and hedonic quality, while still covering the spectrum of product qualities assessed by the original UEQ.
Cognitive Load Scale (CLS) The Cognitive Load Scale (CLS) is a ten-item inventory, developed and validated by Leppink et al. (Leppink et al., 2013). The scale is based on the cognitive load theory (Van Merrienboer and Sweller, 2005) which posits that instructions or instructional material can impose three types of cognitive load on the learner: intrinsic load (IL), extraneous load (EL), and germane load (GL). Intrinsic load refers to the inherent difficulty of the instructional content, which is influenced by the learners’ prior knowledge. Extraneous load is the unnecessary cognitive effort caused by poorly designed instruction. Germane load, introduced later in the theory, refers to the mental effort consciously invested by learners in processing the intrinsic load (Hadie and Yusoff, 2016).
Epistemically-Related Emotion Scales (EES) The Epistemically-Related Emotion Scales (EES), developed by Pekrun et al., measure multiple emotions during knowledge-generating or epistemic activities. The instrument constists of 7 three-item scales, which assess the emotions of surprise, curiosity, enjoyment, confusion, anxiety, frustration, and boredom.
Feedback Perceptions Questionnaire (FPQ) Developed by (Strijbos et al., 2010), the Feedback Perceptions Questionnaire (FPQ) measures feedback perceptions, in educational contexts, in terms of perceived fairness, usefulness, acceptance, willingness to improve, and affect.
Questionnaire assessing Chinese secondary school students’ intention to learn AI Chai et al. draw on the Theory of planned behavior (Ajzen, 1991) and the TAM (Davis, 1989) literature to identify factors that may affect secondary students’ intention to learn AI. The survey involves nine factors: students’ knowledge about AI, general AI anxiety, and subjective norms are presented as background factors that influence their attitudes toward the behavior (i.e., perceived usefulness, social good, and attitude toward use) as well as perceived behavioral control (confidence and optimism) and their behavioral intention to learn AI.
Attitude Toward Cheating (ATC) scale The 34-item ATC scale developed by Gardner and Melvin predicts cheating behavior. Fourteen of the 34 items explicitly condemned cheating, while the remaining 20 items expressed tolerance toward cheaters. Since cheating is generally regarded negatively, students may be inclined to hide their true feelings. To promote objectivity and honesty, the items were worded without referencing the readers; that is, no items included personal pronouns. This approach ensured that the readers were not put on the defensive regarding their own behavior; instead, their attitudes were inferred from their judgments of others’ behavior.
Pupils’Attitudes Towards Technology Short Questionnaire (PATT) Pupils’Attitudes Towards Technology (PATT) was first developed by de1988techniek and has been recently reconstructed and revalidated by Ardies et al.. The revision led to the creation of the PATT Short Questionnaire (PATT-SQ), where the original 58 items were reduced to 24, divided into six subfactors. These six subfactors are Career Aspirations, Interest in Technology, Tediousness of Technology, Positive Perception of Effects of Technology, Perception of Difficulty, and Perception of Technology as a Subject for Boys or for Both Boys and Girls. The PATT-SQ was examined and further developed by Svenningsson et al., leading to the creation of the PATT-SQ-SE.
General Attitudes towards AI Scale (GAAIS) The General Attitudes towards AI Scale (GAAIS) was developed based on the rationale that older Technology Acceptance Scales, such as the TAM (Davis, 1989), reflect users’ individual choices to use technology, while AI often involves decisions made by others. The scale comprises 32 items divided into two subscales: one measuring positive attitudes (Positive GAAIS) and the other measuring negative attitudes (Negative GAAIS). The Positive GAAIS consists of questions related to utility, desired use, and emotions (such as excitement), while the Negative GAAIS contains items related to unethical uses, errors, and negative emotions associated with using AI technology.
Attitudes Towards Using ChatGPT (ATUC) The Attitudes Towards Using ChatGPT (ATUC) scale was created by Ajlouni et al. based on the ABC Model of Attitudes to evaluate undergraduate students’ attitudes toward ChatGPT as a learning tool. The ATUC comprises 22 items (6–26) distributed across three subscales: a) the affects subscale (comprising seven items, 6–12), b) the cognitive subscale (comprising eight items, 13–20), and c) the behavioral subscale (comprising seven items, 21–27). Student Attitudes Toward Artificial Intelligence (SATAI) scale The SATAI scale was developed by Suh and Ahn and consists of 26 items that encompass three components—cognitive, affective, and behavioral factors—representing K-12 students’ attitudes toward AI.
Computer Technology Use Scale (CTUS) The Computer Technology Use Scale was developed by Conrad and Munro and includes five factors: computer efficacy, technology-related anxiety, complexity, positive attitudes, and negative attitudes. These factors are organized into three domains: computer efficacy (perceived capacity), attitudes toward technology, and technology-related anxiety.
Technology Acceptance Model Edited to Assess ChatGPT Adoption (TAME-ChatGPT) The Technology Acceptance Model Edited to Assess ChatGPT Adoption (TAME-ChatGPT) survey was created based on the TAM framework (Sallam et al., 2023). It comprised 13 items for participants who heard of ChatGPT but did not use it and 23 items for participants who used ChatGPT.